OpenStudy (anonymous):

Solve the equation. 3p-1=5(p-1)-2(7-2p)

OpenStudy (muzzack):

3

OpenStudy (muzzack):

easy plz medal me :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

3p – 1 = 5(p – 1) – 2(7 – 2p) First expand the brackets; 3p – 1 = 5p – 5 – 14 + 4p Sort like terms; -1 + 5 + 14 = 5p + 4p - 3p 18 = 6p p = 18/6 p = 3